OpenStudy (astrophysics):

Find the area of the region that lies inside the first curve and outside the second curve

OpenStudy (astrophysics):

\[r=3\cos \theta~~~~r = 1+\cos \theta \] Diagram:  @ganeshie8 When you have problems as such does the basic execution of integrals still apply, such as f(x)-g(x) upper - lower?

OpenStudy (astrophysics):

\[A=2 \int\limits_{0}^{\pi/3} 1/2 [(3\cos \theta)^2-(1+\cos \theta)^2] d \theta\] like this would be your integral right? Is it always like this?

ganeshie8 (ganeshie8):

the key thing is to sketch the are we seek and finding the intersection points

ganeshie8 (ganeshie8):

setting up integral becomes easy and makes more sense after we get some hold of how the region looks like

OpenStudy (astrophysics):

Yeah I got that part, just trying to understand if it's always similar as such, also r = r to find theta but I did x2 since it was symmetrical otherwise I'd have to do pi/3 to 5pi/3 as well right? But mainly my question is, are these questions always like this or are there problems where you'd add the x and y?

OpenStudy (astrophysics):

I should be asking why \[A=2 \int\limits\limits_{0}^{\pi/3} 1/2 [(3\cos \theta)^2-(1+\cos \theta)^2] d \theta\] is not \[A=2 \int\limits\limits_{0}^{\pi/3} 1/2 [(3\cos \theta)^2+(1+\cos \theta)^2] d \theta\] sorry if this is confusing..

ganeshie8 (ganeshie8):

I'll answer that question on why we're subtracting, but let me point out that integral is wrong

ganeshie8 (ganeshie8):

your integral gives below area, right ?